OpenStudy (anonymous):

verify cos(sin^- 1 x )=sqrt(1-x^2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how are you supposed to 'verify' it? a picture will do

OpenStudy (anonymous):

i am supposed to type out each step. also i can only mess with one side

OpenStudy (anonymous):

ok \(\sin^{-1}(x)=\theta\iff \theta =\sin(x)\) then since \(\sin^2(\theta)+\cos^2(\theta)=1\) you get \[\cos^2(\theta)=1-\sin^2(\theta)=1-x^2\]

OpenStudy (anonymous):

or if you prefer you can start with \[\cos^2(\theta)=1-\sin^2(\theta)\] and so \[\cos^2(\sin^{-1}(x))=1-\sin^2(\sin^{-1}(x))=1-x^2\]

OpenStudy (anonymous):

thank you!!!